# Circuits and Backdoors: Five Shades of the SETH

### v1

@@ 1.2 Summary of Results
Meta-complexity.
Is is worth stressing that, even though we present Figure 1 as a collection of “complexity classes”, and indeed we draw a parallel with classical such classes, the results we present are in a sense orthogonal to normal notions of time complexity. What we mean by this is that, as can be seen in Figure 1, we often show equivalence for problems which (are widely believed to) belong in distinct classes, namely, problems which are FPT on the one hand and W[SAT]- or W[P]-complete on the other. In other words, the equivalence classes of Figure 1 do not indicate that all the problems in each class have the same time complexity; but rather that if we could improve the running time of the best algorithm for one, we could improve it for the others, albeit obtaining different types of running times for each problem. We remark that a similar situation holds for the pw-SETH [31].
@@ 1.3 Overview of Techniques
    For the hypotheses which turn out to be equivalent to the LD-C-SETH, one of the main technical tools we rely on is Barrington’s theorem [5], which allows us to transform a circuit on $n$ inputs and depth $\epsilon n$ into a branching program of constant width and length exponential in $\epsilon n$. This program can then be simulated by a formula with constant pathwidth and a modulator representing the inputs. For the converse direction we need to show that an algorithm disproving the LD-C-SETH can be used to decide the satisfiability of a CNF $\phi$ with a modulator $M$ of size $m$, such that deleting $M$ produces a formula of tree-depth at most $O(\log|\phi|)$. For this, we need to eliminate all the variables of the formula except those of $M$. We show how to do this in a way that is exponential in the tree-depth (and hence polynomial in the size of the formula).
@@
[…]
@@ 1.4 Previous Work
As mentioned, SETH-based lower bounds are an area of very active research in parameterized complexity. Beyond the results following [33], which establish running time lower bounds of the form $(c-\varepsilon)^{k}n^{O(1)}$, for some parameter $k$ and fixed constant $c$, the SETH has also been used to obtain lower bounds for non-FPT problems. The most famous example is perhaps the result of Patrascu and Williams that $k$-Dominating Set cannot be solved in time $n^{k-\varepsilon}$, under the SETH [40].
@@ 1.5 Discussion and Directions for Further Work
For some more concrete research directions, we mention that in the interest of keeping the presentation manageable we have only given a limited number of examples for each equivalence class. It would of course be interesting to discover some more such problems. In particular, one obvious question is to consider the remaining lower bounds given in [15], especially their lower bound for Independent Set parameterized by $(\sigma,\delta)$-hub. Breaking this bound is shown to be at least as hard as the SETH in [15], while conversely the MaxSat-SETH can be shown to be at least as hard, via the equivalence we show with the $(\sigma,\delta)H$-MaxSat-SETH. In other words, Independent Set probably belongs in one of the two bottom classes of Figure 1 for parameter $(\sigma,\delta)$-hub, so it would be interesting to determine which one. Another question to consider is finding a natural characterization for SAT parameterized by a modulator to logarithmic treewidth, as we have already considered constant treewidth, pathwidth, and tree-depth, as well as logarithmic pathwidth and tree-depth.
@@ 1.6 Paper Organization
We first remark that we have tried to use a self-explanatory naming scheme, where X-SETH means the hypothesis that it is impossible to improve upon brute-force search for SAT parameterized by a parameter X, or when X is a complexity class, for the complete problem of X. Nevertheless, the definitions of all hypotheses used in this paper (and shown in Figure 1) and the most basic relations between them are given in Appendix A. Even though Appendix A serves as a useful reference, we have made an effort to give precise statements in all theorems, making sure that potentially non-trivial details in the formulations of various hypothesis, such as the order of quantifications of some constants, are clearly spelled out. For the rest of the paper, after some preliminaries, we present the basic connections for the two bottom classes of Figure 1 in Section 3, and then present one class in each of Section 4, Section 5, and Section 6. In Section 7 we present one further characteristic example for each class.
@@ 2 Preliminaries
We assume the reader is familiar with the basics of parameterized complexity as given for example in [12]. We refer the reader there for the definitions of treewidth and pathwidth. A graph $G$ has tree-depth $k$ if there is a bijective mapping of the vertices of $G$ to the vertices of a rooted tree of depth $k$ so that for every edge $uv$, the images of $u,v$ have an ancestor-descendant relation. Following [15], for integers $\sigma,\delta$, a $(\sigma,\delta)$-hub of a graph $G$ is a set of vertices $M$ such that each component of $G-M$ has size at most $\sigma$ and at most $\delta$ neighbors in $M$.
A modulator of a graph is a set of vertices such that removing it leaves a graph with a desired property. When $G$ is a graph and $M$ a modulator we will often write $G-M$ to denote the graph obtained after removing the modulator.
@@
We will also consider satisfiability problems for general formulas and circuits. A circuit is a DAG where vertices of positive in-degree are labeled as disjunctions ($\lor$), conjunctions ($\land$), or negations ($\neg$) and there is a unique sink, called the output gate. Vertices of zero in-degree are called input gates and the satisfiability problem asks if there is a truth assignment to the inputs that makes the output gate evaluate to True. A formula is a circuit where the maximum out-degree of any non-input gate is $1$. Unless otherwise stated we will assume that the fan-in (in-degree) of all circuits is bounded, but in most cases unbounded fan-in gates can be replaced by trees of bounded fan-in gates. The depth of a circuit is the largest distance from an input gate to the output. The size of a circuit is the number of arcs of the corresponding DAG.
Given a CNF formula $\phi$, a strong backdoor set is a set of variables such that for each assignment to these variables we obtain a formula that belongs to a base class. We will only deal with strong backdoors, so we will simply refer to the as backdoors. Furthermore, we will deal with 2-SAT and Horn backdoors, where a formula is Horn if each clause contains at most a positive literal. Recall that 2-SAT is NL-complete and Horn-SAT is P-complete, therefore both problems are solvable in polynomial time. This means that for a given backdoor the satisfiability of $\phi$ can be decided in time $2^{b}|\phi|^{O(1)}$ by trying all assignments to the backdoor. The term backdoor was introduced in [45]. Note that backdoors and modulators are closely related concepts and indeed the notion of modulator to constant treewidth which we use in this paper has also been called a backdoor to constant treewidth [20]. For more information on backdoors see [19].
@@
We give a detailed list of the complexity hypotheses that appear in this paper as well as simple observations about their relations in Appendix A.
@@ 3 Problems Equivalent to (Max-)CNF-SAT
In this paper we offer a more nuanced view that clarifies a bit more the picture presented in [15]. In particular, we show that for some of their lower bound results, not only is the SETH not too strong, but actually the lower bounds they obtain are equivalent to the SETH. We achieve this mainly via Theorem 3.1 below, which shows that the SETH is equivalent to the tdM-SETH, that is the hypothesis that it is impossible to avoid a brute-force enumeration for a given modulator to a graph of bounded tree-depth. Since $(\sigma,\delta)$-hub is a parameter that is a restriction of modulators to bounded tree-depth, we obtain that avoiding brute force enumeration for this parameter is also equivalent to the SETH (Corollary 3.3). In other words, Theorem 3.1 collapses three seemingly distinct parameters (the number of variables, the size of a $(\sigma,\delta)$-hub, and the size of a constant tree-depth modulator) into a single hypothesis. As we show in Section 7, these results establish that the lower bound of [15] for Coloring is actually equivalent to the SETH, and this remains true even if we parameterize by tree-depth modulator rather than $(\sigma,\delta)$-hub.
@@
Despite the above, our results do not indicate that the SETH is the “correct” assumption for all the lower bounds of [15]. In the second part of this section we deal with the MaxSat-SETH, the hypothesis stating the same bound as the SETH but for the problem of maximizing the number of satisfied clauses of a CNF formula. Since this is clearly a problem at least as hard as deciding if a CNF formula can be satisfied, this assumption is at least as plausible as the SETH. Indeed, this assumption is also implied by the so-called Max-3Sat-Hypothesis, which states the same time bound but for CNFs of maximum arity $3$. As mentioned in [15], the relation between the SETH and the Max-3Sat-Hypothesis is currently unknown, which can perhaps be seen as a reason to believe the MaxSat-SETH as more plausible than either of the other two hypothesis. Our second result in this section is that the MaxSat-SETH remains equivalent if we parameterize not by the number of variables but by $(\sigma,\delta)$-hub size (Theorem 3.5). This allows us to establish in Section 7 that lower bounds concerning optimization problems, such as the one on Max Cut parameterized by $(\sigma,\delta)$-hub in [15], are more likely to be equivalent to MaxSat-SETH, hence perhaps more believable, than lower bounds equivalent only to the SETH.
@@
Before we go on, let us mention that the results of this section rely on a precise formulation of the SETH (and related hypotheses) which assumes hardness for $k$-CNF formulas, for sufficiently large $k$. We note that this is important in this section, because we are dealing with extremely restrictive graph parameters. Once we move on to Section 4, it will immediately become possible to break down large clauses into smaller ones by introducing new variables, so the arity of the considered CNFs will no longer be of material importance (see Section A.2).
@@ 3.1 Equivalent Formulations of the SETH
The second statement clearly implies the first, as the set of all variables is a modulator to tree-depth 00. The interesting direction is to show that if we have a $k$-SAT algorithm running in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$, then this implies a faster algorithm parameterized by constant tree-depth modulator.
@@
Recall that resolution is the following rule: when a CNF formula contains two clauses $C,C^{\prime}$ such that for some variable $x$, $C$ contains $x$ and $C^{\prime}$ contains $\neg x$, then we can add to the formula the clause the contains all literals of $C\cup C^{\prime}$ except $x,\neg x$ without affecting satisfiability. We will say that we exhaustively resolve over variable $x$ if we (i) apply this rule to each pair of clauses $C,C^{\prime}$ containing $x,\neg x$ (ii) after doing so remove from $\phi$ all clauses that contain $x$ or $\neg x$.
@@
It is easy to see that the clauses added to $\phi$ do not affect satisfiability, as they can be inferred from existing clauses. To see that deleting all clauses that contain the variable $x$ does not affect satisfiability, we take two cases. If the new formula has a satisfying assignment $\sigma$ such that for all clauses $C$ which contain $x$ positive in $\phi$, a literal of $C$ other than $x$ is set to True by $\sigma$, then we can satisfy $\phi$ by extending $\sigma$ by setting $x$ to False. Otherwise, we claim that all clauses $C^{\prime}$ that contain $\neg x$ in $\phi$ have a literal other than $\neg x$ set to True by $\sigma$, so we can set $x$ to False and satisfy $\phi$. Indeed, we are now in a case where some clause $C$ which contains $x$ positive has all its literals set to False by $\sigma$, but this clause has been paired with all $C^{\prime}$ which contained $\neg x$. Since the new formula is satisfied by $\sigma$, each $C^{\prime}$ must contain a True literal.
@@
We now proceed with the inductive step. Suppose that the claim is true for smaller values of $c$ and we have an instance of maximum arity $k$ and $m$ clauses. By the same arguments as before, applying exhaustive resolution to all leaves produces an instance where $\phi-M$ has tree-depth $c-1$, the maximum arity is at most $k^{\prime}\leq 2k$ and the number of clauses is at most $m^{\prime}=m^{2}$. By inductive hypothesis, completing the procedure will produce a formula with arity at most $2^{c-1}k^{\prime}\leq 2^{c}k$ and with at most $(m^{\prime})^{2^{c-1}}\leq m^{2^{c}}$ clauses, as desired.
@@
The proof consists of observing that a $(\sigma,\delta)$-hub is a modulator to tree-depth at most $\sigma$ and invoking Theorem 3.1.

### v2

@@ 1.2 Summary of Results
Relation with Time Complexity Classes.
It is worth stressing that, even though we present Figure˜1 as a collection of “complexity classes”, and indeed we draw a parallel with classical such classes, the results we present are in a sense orthogonal to normal notions of time complexity. What we mean by this is that, as can be seen in Figure˜1, we often show equivalence for problems which (are widely believed to) belong in distinct classes, namely, problems which are FPT on the one hand and W[SAT]- or W[P]-complete on the other. In other words, the equivalence classes of Figure˜1 do not indicate that all the problems in each class have the same time complexity; but rather that if we could improve the running time of the best algorithm for one, we could improve it for the others, albeit obtaining different types of running times for each problem. We remark that a similar situation holds for the pw-SETH [36].
logpw-SETHLD-C-SETHMax-SAT-SETHSETHWIS Interval+kvWIS Cograph+kvWIS Block+kvIS Cograph+kvIS Block+kvTheorem˜8.18BinWIS Cluster+kvWIS Cluster+kvIS Cluster+kvTheorem˜8.10Theorem˜8.12Theorem˜8.16Theorem˜8.14Theorem˜8.8Theorem˜8.6[19]
Figure 2: Summary of our results for (Weighted) Independent Set parameterized by vertex deletion distance to four graph classes: block, cluster, interval graphs, and cographs. Arrows indicate implication and a problem name indicates the hypothesis that brute-force is inevitable for this case. We consider unweighted versions (IS), unary weighted (WIS) and binary weighted (BinWIS).
A showcase application.
At this point we have established a framework based on five version of the SETH which are, we believe, quite natural and roughly corresponds to five complexity classes, endowed with nice complete problems. Nevertheless, it needs to be said that complexity classes can only be considered as interesting as the problems they contain or capture, or as Papadimitriou and Yannakakis put it, classes that contain interesting problems are discovered not invented [45]. Trying to verify that this is the case for our classes, we return to the motivating example this paper began with: (Weighted) Independent Set on $\mathcal{C}+kv$ graphs, where $\mathcal{C}$ are the graph classes we cherry-picked: cluster graphs, cographs, block graphs, and interval graphs. Our results are summarized in Figure˜2. We identify the LD-C-SETH as a dividing line for the complexity of this problem: for $\textrm{Cluster}+kv$ graphs the problem is at most as hard as the LD-C-SETH, even if binary weights are allowed, while for $\textrm{Cograph}+kv$ and $\textrm{Block}+kv$ the problem is exactly as hard as the LD-C-SETH, in both the unweighted case and the weighted case (with unary weights). However, Weighted Independent Set (again with unary weights) appears to be strictly harder for $\textrm{Interval}+kv$ graphs, as we show that this version is equivalent to the $\log\!\textrm{pw}$M-SETH. Observe that in this way the framework allows us to make comparisons which are not obvious graph-theoretically: cographs, block graphs, and interval graphs are incomparable classes, so a priori one would not have necessarily expected cographs and block graphs to be equivalent but interval graphs to be harder. Of course, a broader advantage is that integrating these questions in our framework allows us to compare these questions with many other, seemingly unrelated questions (such as whether $q$-Coloring can be solved faster than brute-force on graphs with a small modulator to constant pathwidth, which we discussed previously). We believe these results show that this is a direction that merits further investigation.
@@ 1.3 Overview of Techniques
    For the hypotheses which turn out to be equivalent to the LD-C-SETH, one of the main technical tools we rely on is Barrington’s theorem [6], which allows us to transform a circuit on $n$ inputs and depth $\varepsilon n$ into a branching program of constant width and length exponential in $\varepsilon n$. This program can then be simulated by a formula with constant pathwidth and a modulator representing the inputs. For the converse direction we need to show that an algorithm disproving the LD-C-SETH can be used to decide the satisfiability of a CNF $\phi$ with a modulator $M$ of size $m$, such that deleting $M$ produces a formula of tree-depth at most $O(\log|\phi|)$. For this, we need to eliminate all the variables of the formula except those of $M$. We show how to do this in a way that is exponential in the tree-depth (and hence polynomial in the size of the formula). A particular technical challenge here is that, to generalize this result to MaxSAT (which will be crucial for later applications) we need to extend these arguments to arithmetic circuits and then move back to boolean circuits to obtain equivalence with the LD-C-SETH.
@@
    For the equivalence between 2SatB-SETH and $\log\!\textrm{pw}$M-SETH we rely on the intuition that both 2-SAT and satisfiability for SAT formulas of pathwidth $O(\log|\phi|)$ are essentially reachability problems. However, a closer look reveals that a direct reduction is non-obvious, as deciding 2-SAT is in fact more readily seen to be equivalent to non-reachability. We therefore need to transfer the classical Immerman–Szelepcsényi theorem, which established that NL=coNL, to our setting [29].
@@
  5. 5.
    Finally, for our applications to Independent Set on graph classes, we need to rely on several unrelated ideas. In particular, we use standard but non-obvious facts from circuit complexity, such as the fact that $n$ $n$-bit integers can be summed by a bounded fan-in circuit of depth $O(\log n)$; we use the fact that the circuits of the LD-C-SETH can be considered acyclic, which in turn allows us to encode them using the tree structure hidden in cographs and block graphs; we use the fact that the natural DP algorithm for Weighted Independent Set on interval graphs can be translated to a reachability question to obtain equivalence with the $\log\!\textrm{pw}$M-SETH in this case.
@@ 1.4 Previous Work
[…]
@@ 1.5 Discussion and Directions for Further Work
More broadly, an interesting research direction would of course be to further populate the classes introduced in this paper. Specific questions include:
  1. 1.
    Characterize the complexity of Independent Set parameterized by a modulator to an unweighted interval graph. In the weighted case, the problem is equivalent to the $\log\!\textrm{pw}$M-SETH; could it be easier (for instance, equivalent to the LD-C-SETH) for unweighted graphs?
  2. 2.
    As mentioned, Jaffke and Jansen [32] showed that brute force is unavoidable for $q$-Coloring parameterized by a modulator to a path-forest. Our result improves upon this in one sense (because we show equivalence to the LD-C-SETH, rather than SETH-hardness), but is weaker in another (because we parameterize by a modulator to a graph of constant pathwidth, rather than a path). Can this gap be closed? We conjecture that $q$-Coloring parameterized by feedback vertex set or modulator size to a path-forest is in fact easier than the LD-C-SETH. Evidence for such a claim could be established, for instance, by coming up with a reduction from this problem to CNF-SAT.
  3. 3.
    In a similar vein, a natural question to ask is the precise complexity of Independent Set parameterized by feedback vertex set. Again, our results can probably be used to show hardness (indeed LD-C-SETH-equivalence) when we have a modulator to a graph of constant treewidth, but not for a graph of treewidth $1$. What is the smallest treewidth $c$ for which parameterizing Independent Set by vertex-deletion distance to treewidth $c$ is LD-C-SETH-equivalent?
@@ 1.6 Paper Organization
We first remark that we have tried to use a self-explanatory naming scheme, where X-SETH means the hypothesis that it is impossible to improve upon brute-force search for SAT parameterized by a parameter X, or when X is a complexity class, for the complete problem of X. Nevertheless, the definitions of all hypotheses used in this paper (and shown in Figure˜1) and the most basic relations between them are given in Appendix˜A. Even though Appendix˜A serves as a useful reference, we have made an effort to give precise statements in all theorems, making sure that potentially non-trivial details in the formulations of various hypotheses, such as the order of quantifications of some constants, are clearly spelled out. For the rest of the paper, after some preliminaries, we present the basic connections for the two bottom classes of Figure˜1 in Section˜3, and then present one class in each of Section˜4, Section˜5, and Section˜6. In Section˜7 we present one further characteristic example for each class. Finally, in Section˜8 we present our results on Independent Set summarized in Figure˜2.
@@ 2 Preliminaries
We assume the reader is familiar with the basics of parameterized complexity as given for example in [16]. We refer the reader there for the definitions of treewidth and pathwidth. A graph $G$ has tree-depth $k$ if there is a bijective mapping of the vertices of $G$ to the vertices of a rooted tree of depth $k$ so that for every edge $uv$, the images of $u,v$ have an ancestor-descendant relation. We sometimes call such a mapping a tree-depth decomposition. Following [19], for integers $\sigma,\delta$, a $(\sigma,\delta)$-hub of a graph $G$ is a set of vertices $M$ such that each component of GMG-M has size at most $\sigma$ and at most $\delta$ neighbors in $M$.
A modulator of a graph is a set of vertices such that removing it leaves a graph with a desired property. When $G$ is a graph and $M$ a modulator we will often write GMG-M to denote the graph obtained after removing the modulator.
@@
We will also consider satisfiability problems for general formulas and circuits. A boolean circuit is a DAG where vertices of positive in-degree are labeled as disjunctions ($\lor$), conjunctions ($\land$), or negations ($\neg$) and there is a unique sink, called the output gate. Vertices of zero in-degree are called input gates and the satisfiability problem asks if there is a truth assignment to the inputs that makes the output gate evaluate to True. A formula is a circuit where the maximum out-degree of any non-input gate is $1$. Unless otherwise stated we will assume that the fan-in (in-degree) of all circuits is bounded, but in most cases unbounded fan-in gates can be replaced by trees of bounded fan-in gates. The depth of a circuit is the largest distance from an input gate to the output. The size of a circuit is the number of arcs of the corresponding DAG. We will also consider arithmetic circuits, where the value of each gate is an integer rather than a boolean and will mainly be concerned with circuits where the gates implement the addition (++) and maximum ($\max$) operations, that is, they output the sum or maximum of their inputs respectively. Boolean circuits can be interpreted as arithmetic circuits via the natural correspondence that False is 0 and True is $1$. To make the converse transition, we will make use of the following known facts (see e.g. [11]).
Theorem 2.1.
Given an integer $n$ it is possible to construct in time polynomial in $n$ bounded fan-in boolean circuits with $n^{2}$ inputs, $O(n)$ outputs, size $n^{O(1)}$ and depth $O(\log n)$ solving the following problems: (i) computing the maximum of $n$ integers of $n$ bits each (ii) computing the sum of $n$ integers of $n$ bits each. Furthermore, there exists a bounded fan-in boolean circuit with $2n$ inputs, encoding two $n$-bit binary numbers, that has depth $O(\log n)$ and decides if the first number is larger than the second.
Using Theorem˜2.1 we will be able to convert an arithmetic circuit which only uses $\max$ and ++ gates into a boolean circuit, without increasing the depth by much, even if the $\max$ and ++ gates do not have bounded fan-in (because Theorem˜2.1 talks about computing the sum or maximum of $n$ inputs).
Given a CNF formula $\phi$, a strong backdoor set is a set of variables such that for each assignment to these variables we obtain a formula that belongs to a base class. We will only deal with strong backdoors, so we will simply refer to them as backdoors. Furthermore, we will deal with 2-SAT and Horn backdoors, where a formula is Horn if each clause contains at most one positive literal. Recall that 2-SAT is NL-complete and Horn-SAT is P-complete, therefore both problems are solvable in polynomial time. This means that for a given backdoor the satisfiability of $\phi$ can be decided in time $2^{b}|\phi|^{O(1)}$ by trying all assignments to the backdoor. The term backdoor was introduced in [51]. Note that backdoors and modulators are closely related concepts and indeed the notion of modulator to constant treewidth which we use in this paper has also been called a backdoor to constant treewidth [24]. For more information on backdoors see [23].
@@
We give a detailed list of the complexity hypotheses that appear in this paper as well as simple observations about their relations in Appendix˜A.
@@ 3 Problems Equivalent to (Max-)CNF-SAT
In this paper we offer a more nuanced view that clarifies a bit more the picture presented in [19]. In particular, we show that for some of their lower bound results, not only is the SETH not too strong, but actually the lower bounds they obtain are equivalent to the SETH. We achieve this mainly via Theorem˜3.1 below, which shows that the SETH is equivalent to the tdM-SETH, that is the hypothesis that it is impossible to avoid a brute-force enumeration for a given modulator to a graph of bounded tree-depth. Since $(\sigma,\delta)$-hub is a parameter that is a restriction of modulators to bounded tree-depth, we obtain that avoiding brute force enumeration for this parameter is also equivalent to the SETH (Corollary˜3.3). In other words, Theorem˜3.1 collapses three seemingly distinct parameters (the number of variables, the size of a $(\sigma,\delta)$-hub, and the size of a constant tree-depth modulator) into a single hypothesis. As we show in Section˜7, these results establish that the lower bound of [19] for Coloring is actually equivalent to the SETH, and this remains true even if we parameterize by tree-depth modulator rather than $(\sigma,\delta)$-hub.
@@
Despite the above, our results do not indicate that the SETH is the “correct” assumption for all the lower bounds of [19]. In the second part of this section we deal with the MaxSat-SETH, the hypothesis stating the same bound as the SETH but for the problem of maximizing the number of satisfied clauses of a CNF formula. Since this is clearly a problem at least as hard as deciding if a CNF formula can be satisfied, this assumption is at least as plausible as the SETH. Indeed, this assumption is also implied by the so-called Max-3Sat-Hypothesis, which states the same time bound but for CNFs of maximum arity $3$. As mentioned in [19], the relation between the SETH and the Max-3Sat-Hypothesis is currently unknown, which can perhaps be seen as a reason to believe the MaxSat-SETH as more plausible than either of the other two hypothesis. Our second result in this section is that the MaxSat-SETH remains equivalent if we parameterize not by the number of variables but by $(\sigma,\delta)$-hub size (Theorem˜3.5). This allows us to establish in Section˜7 that lower bounds concerning optimization problems, such as the one on Max Cut parameterized by $(\sigma,\delta)$-hub in [19], are more likely to be equivalent to MaxSat-SETH, hence perhaps more believable, than lower bounds equivalent only to the SETH.
@@
Before we go on, let us mention that the results of this section rely on a precise formulation of the SETH (and related hypotheses) which assumes hardness for $k$-CNF formulas, for sufficiently large $k$. We note that this is important in this section, because we are dealing with extremely restrictive graph parameters. Once we move on to Section˜4, it will immediately become possible to break down large clauses into smaller ones by introducing new variables, so the arity of the considered CNFs will no longer be of material importance (see Section˜A.2).
@@ 3.1 Equivalent Formulations of the SETH
The second statement clearly implies the first, as the set of all variables is a modulator to tree-depth 0. The interesting direction is to show that if we have a $k$-SAT algorithm running in time $(2-\varepsilon)^{n}|\phi|^{O(1)}$, then this implies a faster algorithm parameterized by constant tree-depth modulator.
@@
Recall that resolution is the following rule: when a CNF formula contains two clauses $C,C^{\prime}$ such that for some variable $x$, $C$ contains $x$ and $C^{\prime}$ contains $\neg x$, then we can add to the formula the clause that contains all literals of $C\cup C^{\prime}$ except $x,\neg x$ without affecting satisfiability. We will say that we exhaustively resolve over variable $x$ if we (i) apply this rule to each pair of clauses $C,C^{\prime}$ containing $x,\neg x$ (ii) after doing so remove from $\phi$ all clauses that contain $x$ or $\neg x$.
@@
It is easy to see that the clauses added to $\phi$ do not affect satisfiability, as they can be inferred from existing clauses. To see that deleting all clauses that contain the variable $x$ does not affect satisfiability, we take two cases. If the new formula has a satisfying assignment $\sigma$ such that for all clauses $C$ which contain $x$ positive in $\phi$, a literal of $C$ other than $x$ is set to True by $\sigma$, then we can satisfy $\phi$ by extending $\sigma$ by setting $x$ to False. Otherwise, we claim that all clauses $C^{\prime}$ that contain $\neg x$ in $\phi$ have a literal other than $\neg x$ set to True by $\sigma$, so we can set $x$ to True and satisfy $\phi$. Indeed, we are now in a case where some clause $C$ which contains $x$ positive has all its literals set to False by $\sigma$, but this clause has been paired with all $C^{\prime}$ which contained $\neg x$. Since the new formula is satisfied by $\sigma$, each $C^{\prime}$ must contain a True literal.
@@
We now proceed with the inductive step. Suppose that the claim is true for smaller values of $c$ and we have an instance of maximum arity $k$ and $m$ clauses. By the same arguments as before, applying exhaustive resolution to all leaves produces an instance where $\phi-M$ has tree-depth c1c-1, the maximum arity is at most $k^{\prime}\leq 2k$ and the number of clauses is at most $m^{\prime}=m^{2}$. By inductive hypothesis, completing the procedure will produce a formula with arity at most $2^{c-1}k^{\prime}\leq 2^{c}k$ and with at most $(m^{\prime})^{2^{c-1}}\leq m^{2^{c}}$ clauses, as desired.
@@
The proof consists of observing that a $(\sigma,\delta)$-hub is a modulator to tree-depth at most $\sigma$ and invoking Theorem˜3.1.